# Segmenting Action-Value Functions Over Time-Scales in SARSA using TD($\Delta$)

### v1

@@ Acknowledgments
$\sum_{z=1}^{Z}\bigg{(}\gamma_{z}^{k_{z}}-\gamma_{z-1}^{k_{z}}\bigg{)}\sum_{u=0}^{z-1}\Delta_{t-1}^{u}+\gamma_{z}^{k_{z}}\Delta_{t-1}^{z}=\sum_{u=0}^{Z-1}\sum_{z=u+1}^{Z}\bigg{(}\gamma_{z}^{k_{z}}-\gamma_{z-1}^{k_{z}}\bigg{)}\Delta_{t-1}^{u}+\sum_{z=1}^{Z}\gamma_{z}^{k_{z}}\Delta_{t-1}^{z}$
$=\sum_{u=0}^{Z-1}\Delta_{t-1}^{u}\bigg{(}\sum_{z=u+1}^{Z}\gamma_{z}^{k_{z}}-\sum_{z=u}^{Z-1}\gamma_{z}^{k_{z+1}}\bigg{)}+\sum_{z=1}^{Z}\gamma_{z}^{k_{z}}\Delta_{t-1}^{z}$
$=\sum_{u=0}^{Z-1}\Delta_{t-1}^{u}\bigg{(}\sum_{z=u+1}^{Z-1}(\gamma_{z}^{k_{z}}-\gamma_{z}^{k_{z+1}})+(\gamma_{Z}^{k_{z}}-\gamma_{u}^{k_{u+1}})\bigg{)}+\sum_{z=1}^{Z}\gamma_{z}^{k_{z}}\Delta_{t-1}^{z}$
$=\sum_{u=0}^{Z-1}\sum_{z=u+1}^{Z-1}(\gamma_{z}^{k_{z}}-\gamma_{z}^{k_{z+1}})\Delta_{t-1}^{u}+\gamma_{Z}^{k_{z}}\sum_{z=0}^{Z}\Delta_{t-1}^{z}+\sum_{z=0}^{Z-1}(\gamma_{z}^{k_{z}}-\gamma_{z}^{k_{z+1}})\Delta_{t-1}^{z}$
$=\sum_{u=0}^{Z-1}\sum_{z=u}^{Z-1}(\gamma_{z}^{k_{z}}-\gamma_{z}^{k_{z+1}})\Delta_{t-1}^{u}+\gamma_{Z}^{k_{z}}\sum_{z=0}^{Z}\Delta_{t-1}^{z}$
$=\sum_{z=0}^{Z-1}(\gamma_{z}^{k_{z}}-\gamma_{z}^{k_{z+1}})\sum_{u=0}^{z}\Delta_{t-1}^{u}+\gamma_{Z}^{k_{z}}\sum_{z=0}^{Z}\Delta_{t-1}^{z}$ (46)
@@
$\sum_{z=0}^{Z}\Delta_{t}^{z}\leq\underbrace{\epsilon\bigg{(}\frac{1-\gamma^{k}}{1-\gamma}\bigg{)}+\epsilon\bigg{(}\sum_{z=0}^{Z-1}\frac{\gamma_{z}^{k_{z+1}}-\gamma^{k_{z}}_{z}}{1-\gamma_{z}}\bigg{)}}_{\text{variance reduction}}+\underbrace{\sum_{z=0}^{Z-1}\bigg{(}\gamma^{k_{z}}_{z}-\gamma^{k_{z+1}}_{z}\bigg{)}\sum_{u=0}^{z}\Delta_{t-1}^{u}+\gamma^{k}\sum_{z=0}^{Z}\Delta_{t-1}^{z}}_{\text{bias introduction}}$ (47)

### v2

@@ Acknowledgments
$\sum_{m=1}^{M}\bigg{(}\eta_{m}^{k_{m}}-\eta_{m-1}^{k_{m}}\bigg{)}\sum_{q=0}^{m-1}\Delta_{n-1}^{q}+\eta_{m}^{k_{m}}\Delta_{n-1}^{m}=\sum_{q=0}^{M-1}\sum_{m=q+1}^{M}\bigg{(}\eta_{m}^{k_{m}}-\eta_{m-1}^{k_{m}}\bigg{)}\Delta_{n-1}^{q}+\sum_{m=1}^{M}\eta_{m}^{k_{m}}\Delta_{n-1}^{m}$
$=\sum_{q=0}^{M-1}\Delta_{n-1}^{q}\bigg{(}\sum_{m=q+1}^{M}\eta_{m}^{k_{m}}-\sum_{m=q}^{M-1}\eta_{m}^{k_{m+1}}\bigg{)}+\sum_{m=1}^{M}\eta_{m}^{k_{m}}\Delta_{n-1}^{m}$
$=\sum_{q=0}^{M-1}\Delta_{n-1}^{q}\bigg{(}\sum_{m=q+1}^{M-1}(\eta_{m}^{k_{m}}-\eta_{m}^{k_{m+1}})+(\eta_{M}^{k_{m}}-\eta_{q}^{k_{q+1}})\bigg{)}+\sum_{m=1}^{M}\eta_{m}^{k_{m}}\Delta_{n-1}^{m}$
$=\sum_{q=0}^{M-1}\sum_{m=q+1}^{M-1}(\eta_{m}^{k_{m}}-\eta_{m}^{k_{m+1}})\Delta_{n-1}^{q}+\eta_{M}^{k_{m}}\sum_{m=0}^{M}\Delta_{n-1}^{m}+\sum_{m=0}^{M-1}(\eta_{m}^{k_{m}}-\eta_{m}^{k_{m+1}})\Delta_{n-1}^{m}$
$=\sum_{q=0}^{M-1}\sum_{m=q}^{M-1}(\eta_{m}^{k_{m}}-\eta_{m}^{k_{m+1}})\Delta_{n-1}^{q}+\eta_{M}^{k_{m}}\sum_{m=0}^{M}\Delta_{n-1}^{m}$
$=\sum_{m=0}^{M-1}(\eta_{m}^{k_{m}}-\eta_{m}^{k_{m+1}})\sum_{q=0}^{m}\Delta_{n-1}^{q}+\eta_{M}^{k_{m}}\sum_{m=0}^{M}\Delta_{n-1}^{m}$ (49)
@@
$\sum_{m=0}^{M}\Delta_{n}^{m}\leq\underbrace{\epsilon\bigg{(}\frac{1-\eta^{k}}{1-\eta}\bigg{)}+\epsilon\bigg{(}\sum_{m=0}^{M-1}\frac{\eta_{m}^{k_{m+1}}-\eta^{k_{m}}_{m}}{1-\eta_{m}}\bigg{)}}_{\text{variance reduction}}+\underbrace{\sum_{m=0}^{M-1}\bigg{(}\eta^{k_{m}}_{m}-\eta^{k_{m+1}}_{m}\bigg{)}\sum_{q=0}^{m}\Delta_{n-1}^{q}+\eta^{k}\sum_{m=0}^{M}\Delta_{n-1}^{m}}_{\text{bias introduction}}$ (50)